OpenStudy (anonymous):

Harold, the happy concrete guy, got a call on the phone about some concrete he has to pour for an odd shaped triangular patio. The only information he got was that the side lengths were 5, 6, and 8 meters long. a.) Find the angles for Harold so he can stay happy. b.) Find the area of the patio that Harold will be forming. Any help? I'm horrible with triangles.

OpenStudy (anonymous):

Hold on a second while I do the calculations

OpenStudy (anonymous):

okay haha thanks

OpenStudy (trancenova):

c2 = a2 + b2 – 2ab cos C so 5^2=6^2+8^2-2(6*8)CosC CosC= (5^2-6^2-8^2)/(-2*6*8) So the angle C = 38.62deg.

OpenStudy (trancenova):

 Area = a*b*sin(C)/2